OpenStudy (anonymous):

√3 - 4√6 + √48 - √54

OpenStudy (anonymous):

\[\sqrt{3}-4\sqrt{6}+\sqrt{16*3}-\sqrt{9*6}\]

OpenStudy (anonymous):

\[\sqrt{3}-4\sqrt{6}+4\sqrt{3}-3\sqrt{6}\]

OpenStudy (jiteshmeghwal9):

\[\sqrt{3}+4\sqrt{6}+4\sqrt{3}-3\sqrt{6}\]now add all the similar term as

OpenStudy (jiteshmeghwal9):

\[5\sqrt{3}+\sqrt{6}\]

OpenStudy (anonymous):

\[5\sqrt{3}-7\sqrt{6}\]

OpenStudy (anonymous):

no those arent one of the multiple choices

OpenStudy (anonymous):

@JohnHanShanghai and @jiteshmeghwal9 those arent even on the multiple choice answers

OpenStudy (anonymous):

here is a calculator

OpenStudy (anonymous):

there answers are right

OpenStudy (anonymous):

@JohnHanShanghai has the right answer

OpenStudy (anonymous):

johnis correct

OpenStudy (anonymous):

but its not on the mutltiple choice

OpenStudy (anonymous):

what are the multiple choices

OpenStudy (anonymous):

@GhettoBaby did u look at the link

OpenStudy (anonymous):

yes and nothing on their either

OpenStudy (anonymous):

the decimal approximation is -8.48617416163786021974375681541187790904760609554478

OpenStudy (anonymous):

rounded off its 8.50

OpenStudy (anonymous):

sorry i wrote the question wrong its 7√3 - 4√6 + √48 - √54

OpenStudy (anonymous):

thank you lol

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

Don't use wolfram for these type of problem. You should be able to solve it by hand

OpenStudy (anonymous):

im in the 8th grade going to high school. don't judge me